Radovilsky Manufacturing​ Company, in​ Hayward, California, makes flashing lights for toys. The company operates its production facility 300 days per year. It has orders for about 11,800 flashing lights per year and has the capability of producing 105 per day. Setting up the light production costs $49. The cost of each light is $1.05. The holding cost is ​$0.10 per light per year.

A. What is the optimal size of the production​ run? ____ units

B.) What is the average holding cost per​ year?

C) What is the average setup cost per​ year?

D. What is the total cost per​ year, including the cost of the​ lights?
Answer #1

A. Economic Production Quantity = Sqroot{(2*Annual Demand*setup Cost)/Holding cost}

= Sqroot{(2*11800*49)/0.1}

= 3400.59

B. Average Holding cost = Average Inventory * Holding cost per unit

= (3400.59/2)*0.1

= $170.03

C. Setup cost = Number of production runs * Setup cost per run

= (11800/3400.59)*49

= $170.03

D. Total cost of lights = Annual Demand * cost of each light

= 11800*1.05

= $12390

Total cost = $12390 + $170.03 + $170.03

= $12730.06